[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Jul 15, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Fri Jul 15 12:43:56 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The rainfall products are not updating. Monica will continue to investigate.

The TAFB shfit log e-mails are not always being sent. TSB will continue to  
investigate.


HSU:
The problem with the ATCF A-deck form Celia was fixed late yesterday thanks  
to some detective work by Monica, Dave, Craig and Mike. The corrupted filed  
needed to be replaced in several places to fix the problem. Monica  
documented the procedure in case this happens again. We still don't know  
the underlying reason why the file got corrupted.

TSB is working with NCO to make the HWRF fields available sooner in  
N-AWIPS. After considering several options, we decided to create the  
N-AWIPS files locally at NHC from the grib files, since the grib files are  
much smaller and are available much sooner from NCO. This development may  
take up to a few weeks.

Storm surge collaboration tests are planned for Mon-Thu of next week using  
the same scenarios as were planned when the test was postponed due to TS  
Colin and technical issues. Mark will make the documentation available, and  
Chris Mello will be available to assist during the tests if needed.


Other items:
Several staff from NCO will be at NHC on Monday. If there are any specific  
issues you would like to bring to their attention let Mark know. We already  
plan to get an update on staffing and the NCO Helpdesk.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160715/58839157/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list